### v1

# Restivo Salemi property for $\alpha$-power free languages with $\alpha\geq 5$ and $k\geq 3$ letters

Abstract
In 2009, Shur published the following conjecture: Let $L$ be a power-free language and let $e(L)\subseteq L$ be the set of words of $L$ that can be extended to a bi-infinite word respecting the given power-freeness. If $u,v\in e(L)$ then $uwv\in e(L)$ for some word $w$.
Let $L_{k,\alpha}$ denote an $\alpha$-power free language over an alphabet with $k$ letters, where $\alpha$ is a positive rational number and $k$ is positive integer. We prove the conjecture for the languages $L_{k,\alpha}$, where $\alpha\geq 5$ and $k\geq 3$.
Keywords: Transition Property; Power Free Words; Restivo Salemi Property
1 Introduction
Let $r$ be a non-empty word, let $\alpha$ be a positive rational constant, and let $r^{\alpha}=rr\cdots rt$ be such that $\frac{|r^{\alpha}|}{|r|}=\alpha$ and $t$ is a prefix of $r$. We call the word $r^{\alpha}$ an $\alpha$-power of $r$. For example $(1234)^{3}=123412341234$, $(1234)^{\frac{7}{4}}=1234123$, and $(1234)^{\frac{1}{2}}=12$.
The square free ($2$-power free) and cube free words ($3$-power free) are well known examples of the power free words. More detailed introduction into the field of power free words can be found, for instance, in [2] and [8].
In 1985, Restivo and Salemi presented a list of five problems concerning the extendability of power free words [3]. Problem $4$ and Problem $5$ are relevant for the current article, hence we recall these two problems:
  • Problem $4$: Given finite $\alpha$-power-free words $u$ and $v$, decide whether there is a transition word $w$, such that $uwu$ is $\alpha$-power free.
  • Problem $5$: Given finite $\alpha$-power-free words $u$ and $v$, find a transition word $w$, if it exists.
In 2019, a survey on the solution of all the five problems has been presented in [1]; specifically Problem $4$ and Problem $5$ were resolved for some binary language. In addition, in [1] the authors resolved Problem $4$ and $5$ for cube free words.
Let $\mathbb{N}$ denote the set of positive integers and let $\mathbb{Q}$ denote the set of rational numbers.
Definition 1.1.
(see [4, Definition $1$])
$\begin{split}\Upsilon=\{(k,\alpha)\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k=3\mbox{ and }\alpha>2\}\\ \cup\{(k,\alpha)\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k>3\mbox{ and }\alpha\geq 2\}\\ \cup\{(k,\alpha^{+})\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k\geq 3\mbox{ and }\alpha\geq 2\}\mbox{.}\end{split}$
Remark 1.2.
(see [4, Remark $1$]) The definition of $\Upsilon$ says that: If $(k,\alpha)\in\Upsilon$ and $\alpha$ is a “number with $+$” then $k\geq 3$ and $\alpha\geq 2$. If $(k,\alpha)\in\Upsilon$ and $\alpha$ is “just” a number then $k=3$ and $\alpha>2$ or $k>3$ and $\alpha\geq 2$.
In 2020, Problem $4$ and Problem $5$ have were resolved for $\alpha$-power free languages over an alphabet with $k$ letters, where $(k,\alpha)\in\Upsilon$ [4].
In 2009, a conjecture related to Problems $4$ and Problem $5$ of Restivo and Salemi appeared in [7]:
Conjecture 1.3.
Let $L$ be a power-free language and let $e(L)\subseteq L$ be the set of words of $L$ that can be extended to a bi-infinite word respecting the given power-freeness. If $u,v\in e(L)$ then $uwv\in e(L)$ for some word $w$.
In 2018, Conjecture 1.3 appeared also in [6] using a “Restivo Salemi property”; it was defined that a language $L$ has the Restivo Salemi property if Conjecture 1.3 holds for the language $L$.
Let $L_{k,\alpha}$ denote an $\alpha$-power free language over an alphabet with $k$ letters, where $\alpha$ is a positive rational number and $k$ is positive integer.
In [6, Remark $41$], it was mentioned that for $2^{+}\leq\alpha\leq\frac{7}{3}$ the language $L_{2,\alpha}$ has the Restivo Salemi property (Conjecture 1.3 holds for $L_{2,\alpha}$), where $2^{+}$-power free language denotes the overlap free language. Otherwise, as to our best knowledge, Conjecture 1.3 remains open.
Let
$\widetilde{\Upsilon}=\{(k,\alpha)\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k\geq 3\mbox{ and }\alpha\geq 5\}\mbox{.}$
Clearly $\widetilde{\Upsilon}\subset\Upsilon$. In the current article we prove that if $(k,\alpha)\in\widetilde{\Upsilon}$ then Conjecture 1.3 holds for the language $L_{k,\alpha}$.
In order to prove our result we apply the ideas and some results from [4] and [5]. We explain briefly the very basic idea. If $u,v$ are $\alpha$-power free words and $x$ is a letter such that $x$ is a factor of neither $u$ nor $v$, then clearly $uxv$ is square free ($2$-power free). Just note that there cannot be a factor in $uxv$ which is a square and contains $x$, because $x$ has only one occurrence in $uxv$. In [4], the author considers right and left infinite words containing a given factor and a non-recurrent letter $x$. Due to the non-recurrence of the letter $x$ and some involved observations of recurrent and non-recurrent factors, the author managed to prove the results. In 2023, the idea of right and left infinite power free words with a non-recurrent letter have been extended to bi-infinite $\alpha$-power free words over an alphabet with $k$-letters, where $(\alpha,k)\in\widetilde{\Upsilon}$, [5]. Moreover, in [5], it was predicted, that these bi-infinite power free words with a non-recurrent letter could help to prove Conjecture 1.3.
2 Preliminaries
Due to a significant overlap of notation, we copy and use the section “Preliminaries” from [5].
Let $\Sigma_{k}$ denote an alphabet with $k$ letters. Let $\Sigma_{k}^{+}$ denote the set of all nonempty finite words over $\Sigma_{k}$, let $\epsilon$ denote the empty word, let $\Sigma_{k}^{*}=\Sigma_{k}^{+}\cup\{\epsilon\}$, let $\Sigma_{k}^{\mathbb{N},R}$ denote the set of all right infinite words over $\Sigma_{k}$, let $\Sigma_{k}^{\mathbb{N},L}$ denote the set of all left infinite words over $\Sigma_{k}$, and let $\Sigma_{k}^{\mathbb{Z}}$ denote the set of all bi-infinite words over $\Sigma_{k}$.
Let $\Sigma_{k}^{\infty}=\Sigma_{k}^{\mathbb{N},L}\cup\Sigma_{k}^{\mathbb{N},R}\cup\Sigma_{k}^{\mathbb{Z}}$. We call $w\in\Sigma_{k}^{\infty}$ an infinite word.
Let $\operatorname{occur}(w,t)$ denote the number of all occurrences of the nonempty factor $t\in\Sigma_{k}^{+}$ in the word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$. If $w\in\Sigma_{k}^{\infty}$ and $\operatorname{occur}(w,t)=\infty$, then we call $t$ a recurrent factor in $w$.
Let $F(w)$ denote the set of all finite factors of a finite or infinite word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$. The set $F(w)$ contains the empty word and if $w$ is finite then also $w\in F(w)$. Let $F_{r}(w)\subseteq F(w)$ denote the set of all recurrent nonempty factors of $w\in\Sigma_{k}^{\infty}$.
Let $\operatorname{Prf}(w)\subseteq F(w)$ denote the set of all prefixes of $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\mathbb{N},R}$ and let $\operatorname{Suf}(w)\subseteq F(w)$ denote the set of all suffixes of $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\mathbb{N},L}$. We define that $\epsilon\in\operatorname{Prf}(w)\cap\operatorname{Suf}(w)$ and if $w$ is finite then also $w\in\operatorname{Prf}(w)\cap\operatorname{Suf}(w)$.
[…]
We define the reversal $w^{R}$ of a finite or infinite word $w=\Sigma_{k}^{*}\cup\Sigma_{k}^{\mathbb{N},R}\cup\Sigma_{k}^{\mathbb{N},L}$ as follows:
  • $\epsilon^{R}=\epsilon$.
  • If $w\in\Sigma_{k}^{+}$ and $w=w_{1}w_{2}\dots w_{m}$, where $w_{i}\in\Sigma_{k}$ and $1\leq i\leq m$, then $w^{R}=w_{m}w_{m-1}\dots w_{2}w_{1}$.
  • If $w\in\Sigma_{k}^{\mathbb{N},L}$ and $w=\dots w_{2}w_{1}$, where $w_{i}\in\Sigma_{k}$ and $i\in\mathbb{N}$, then $w^{R}=w_{1}w_{2}\dots\in\Sigma_{k}^{\mathbb{N},R}$.
  • If $w\in\Sigma_{k}^{\mathbb{N},R}$ and $w=w_{1}w_{2}\dots$, where $w_{i}\in\Sigma_{k}$ and $i\in\mathbb{N}$, then $w^{R}=\dots w_{2}w_{1}\in\Sigma_{k}^{\mathbb{N},L}$.
Remark 2.1.
It is obvious that the reverse function preserves the power-freeness and that every factor of an $\alpha$-power free word is also $\alpha$-power free.
The next proposition is a “reformulation” of Corollary $1$ from [4] using only the notation of the current article.
Proposition 2.2.
(reformulation of [4, Corollary $1$]) If $(k,\alpha)\in\Upsilon$, $v\in L_{k,\alpha}^{\mathbb{N},L}$, $z\in\operatorname{Suf}(v)$, $x\in F_{r}(v)\cap\Sigma_{k}$, $s\in L_{k,\alpha}^{\mathbb{N},L}$, and $x\not\in F(s)$ then there is a finite word $u\in\Sigma_{k}^{*}$ such that $z\in\operatorname{Suf}(su)$ and $su\in L_{k,\alpha}^{\mathbb{N},L}$.
The following elementary lemma was shown in [4]. For the reader’s convenience we present the lemma with the proof.
Lemma 2.3.
(reformulation [4, Lemma $2$]) If $k\geq 3$ and $\alpha>2$ then $L_{k-1,\alpha}^{\mathbb{N},R}\not=\emptyset$.
Proof.
Thue Morse words are well known overlap-free ($2^{+}$-power free) right infinite words on two letters [6]. Let $x\in\Sigma_{k}$ and let $t\in L_{k,\alpha}^{\mathbb{N},R}$ be a Thue Morse word on two letters such that $x\not\in F(t)$. Since $k\geq 3$, we have that such $t$ exists. This ends the proof. ∎
Given $t\in\Sigma_{k}^{\mathbb{N},R}$, let
$\Phi(t)=\{s\in\Sigma_{k}^{\mathbb{N},L}\mid F(s)\subseteq F_{r}(t)\}\mbox{.}$
It means that all factors of $s\in\Phi(t)$ are recurrent factors of $t$. It was shown in [4], that $\Phi(t)$ is nonempty:
Lemma 2.4.
[4, Lemma $4$] If $t\in\Sigma_{k}^{\mathbb{N},R}$ then $\Phi(t)\not=\emptyset$.
The main result of [5] can be formulated as follows:
Theorem 2.5.
If $\alpha\geq 5$, $k\geq 3$, $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, then there are $\widetilde{v}\in L_{k,\alpha}^{\mathbb{Z}}$ and a letter $x\in\Sigma$ such that $w\in F(\widetilde{v})$, $x\in F(w)$, and $\operatorname{occur}(w,x)<\infty$.
Remark 2.6.
However, we will not apply Theorem 2.5 directly in the current article, hence we will require some additional properties of the prefix and the suffix of the constructed bi-infinite word. Instead we will apply [5, Theorem $1$] (see Theorem 4.1 in the current article).
3 Bi-infinite $\alpha$-power words
For the rest of the article suppose that $(k,\alpha)\in\widetilde{\Upsilon}$; it means $k\geq 3$ and $\alpha\geq 5$.
We define two technical sets $\Gamma$ and $\Delta$.
Definition 3.1.
([5, Definition $2$]) Let $\Gamma$ be a set of triples such that $(w,\eta,u)\in\Gamma$ if and only if
  • $w\in\Sigma_{k}^{+}$, $\eta,u\in\Sigma_{k}^{*}$, and
  • if $|u|\leq|w|$ then $|\eta|\geq(\alpha+1)\alpha^{|w|-|u|}|w|\mbox{.}$
Remark 3.2.
([5, Remark $4$]) The set $\Gamma$ contains triples of finite words $w,\eta,u$ such that $w$ is nonempty, and if $u$ is shorter than $w$, then the word $\eta$ is “sufficiently” longer than $w$.
Definition 3.3.
([5, Definition $3$]) Let $\Delta$ be a set of $6$-tuples such that $(s,\sigma,w,\eta,x,u)\in\Delta$ if and only if
  1. 1.
    $s\in\Sigma_{k}^{\mathbb{N},L}$, $\sigma,\eta,u\in\Sigma_{k}^{*}$, $w\in\Sigma_{k}^{+}$, $x\in\Sigma_{k}$,
  2. 2.
    $s\sigma w\eta xu\in L_{k,\alpha}^{\mathbb{N},L}$,
  3. 3.
    $(w,\eta,u)\in\Gamma$,
  4. 4.
    $\operatorname{occur}(s\sigma w,w)=1$, and
  5. 5.
    $x\not\in F(s)\cup F(u)$.
4 Transition words
Theorem 4.1.
([5, Theorem $1$]) If $(s,\sigma,w,\eta,x,\epsilon)\in\Delta$, $t\in L_{k,\alpha}^{\mathbb{N},R}$, and $x\not\in F(t)$ then there is $\widehat{\eta}\in\operatorname{Prf}(\eta)$ such that $s\sigma w\widehat{\eta}xt\in L_{k,\alpha}^{\mathbb{Z}}$.
We show how construct an element of $\Delta$ from a bi-infinite $\alpha$-power free word $v$ with a finite factor $w$.
Lemma 4.2.
If $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, and $x\in F(w)\cap\Sigma_{k}$ then there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$, $\sigma,\eta\in\Sigma_{k}^{*}$ such that $(\overline{s},\sigma,w,\eta,x,\epsilon)\in\Delta$.
Proof.
We distinguish following cases:
  1. 1.
    There are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\eta\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\eta x\widehat{v}$ and $(w,\eta,\epsilon)\in\Gamma$.
  2. 2.
    Otherwise. Then it follows from Case 1 that in this case there are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\overline{\eta}\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\overline{\eta}\widehat{v}$ such that $w\not\in F(\overline{v})$, $x\not\in F(\widehat{v})$, and $(w,\overline{\eta},\epsilon)\not\in\Gamma$.
    Let $u\in\operatorname{Prf}(\widehat{v})$ be such that $|u|>|w\overline{\eta}|$. We claim that $\overline{v}w\overline{\eta}ux\in L_{k,\alpha}^{\mathbb{N},L}$. To get a contradiction, suppose that there are $r$ and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(\overline{v}w\overline{\eta}ux)$. Since $x\not\in F(u)$ and $u\not=\epsilon$ we have that $ux\in L_{k,\alpha}$ and $ux\in\operatorname{Suf}(r)$. Since $|u|>|w\overline{\eta}|$ it follows that $w\in F(rr)$. Since $\beta>\alpha\geq 4$ and $w\not\in F(\overline{v})$, we conclude that are not such $r$ and $\beta$. Set $\eta=\overline{\eta}u$. Obviously $(w,\eta,\epsilon)\in\Gamma$, because $(w,\overline{\eta},\epsilon)\in\Gamma$.
In both cases we found $\overline{v}$ and $\eta$ such that $\overline{v}w\eta x\in L_{k,\alpha}^{\mathbb{N},L}$ and $(w,\eta,\epsilon)\in\Gamma$.
  • If $x\not\in F(\overline{v})$ then let $\overline{s}=\overline{v}$. It is easy to verify that $(\overline{s},\sigma,w,\eta,x)\in\Delta$.
  • If $x\in F(\overline{v})$ then let $t$ be a right infinite $\alpha$-power free word on the alphabet $\Sigma_{k}\setminus\{x\}$. Since $(\alpha,k)\in\widetilde{\Upsilon}$, Lemma 2.3 asserts that $t$ exists. Let $s=t^{R}$. Clearly $s$ is $\alpha$-power free and $x\not\in F(s)$.
    Let $z=w\eta x$. Proposition 2.2 implies that there is a finite word $u$ such that $su\in L_{k,\alpha}^{\mathbb{N},L}$ and $z=w\eta x\in\operatorname{Suf}(su)$. Obviously there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma\in\Sigma_{k}^{*}$ such that $(\overline{s},\sigma,w,\eta,x)\in\Delta$. Note that since $x\in F(w)$ and $x\not\in F(\overline{s})$ Property 4 of Definition 3.3 is easy to be asserted by a proper choice of $\sigma$.
This completes the proof.
Suppose two bi-infinite $\alpha$-power free words $s_{1}\widetilde{w}_{1}t_{1}$ and $s_{2}\widetilde{w}_{2}t_{2}$, where $s_{1},s_{2}$ are left infinite words, $t_{1},t_{2}$ are right infinite words, $w_{1},w_{2}$ are finite words, the factors of $s_{2}$ are recurrent factors of $t_{1}$, and there is a letter $x$ such that $x$ is not a factor of $s_{1},s_{2},t_{1},t_{2}$ and $x$ is a factor of $w_{2}$. We show that there is a finite word $w$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}$ is a bi-infinite $\alpha$-power free word.
Proposition 4.3.
If $\widetilde{w}_{1},\widetilde{w}_{2}\in\Sigma_{k}^{+}$, $s_{1},s_{2}\in\Sigma_{k}^{\mathbb{N},L}$, $t_{1},t_{2}\in\Sigma_{k}^{\mathbb{N},R}$, $x\in\Sigma_{k}$, $s_{1}\widetilde{w}_{1}t_{1},s_{2}\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $F(s_{2})\subseteq F_{r}(t_{1})$, $x\not\in F(s_{1})\cup F(s_{2})\cup F(t_{1})\cup F(t_{2})$, and $x\in F(\widetilde{w}_{2})$ then there is $w\in\operatorname{Prf}(t_{1})$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
Proof.
Let $p\in\operatorname{Suf}(s_{2})$ be the shortest suffix such that
$|p|>\max\{|\widetilde{w}_{1}|,|\widetilde{w}_{2}|\}\mbox{.}$ (1)
Let $h\in\operatorname{Prf}(t_{1})$ be the shortest prefix of $t_{1}$ such that $hp\in\operatorname{Prf}(t_{1})$ and $|h|>|p|$; such $h$ exists, because $F(s_{2})\subseteq F_{r}(t_{1})$. We have that $s_{1}\widetilde{w}_{1}hp\in L_{k,\alpha}^{\mathbb{N},L}$, since $hp\in\operatorname{Prf}(t_{1})$ and $s_{1}\widetilde{w}_{1}t_{1}\in L_{k,\alpha}^{\mathbb{Z}}$. We show that $s_{1}\widetilde{w}_{1}hp\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
To get a contradiction, suppose that there are $g\in\operatorname{Prf}(\widetilde{w}_{2}xt_{2})$, $y\in\Sigma_{k}$, $r\in\Sigma_{k}^{+}$, and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(s_{1}\widetilde{w}_{1}hpgy)$, $gy\in\operatorname{Prf}(\widetilde{w}_{2}xt_{2})$, and $s_{1}\widetilde{w}_{1}hpg\in L_{k,\alpha}^{\mathbb{N},L}$. Since $p\widetilde{w}_{2}xt_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $x\in F(w_{2})$ we have that $|r^{\beta}|>|pgy|$ and $x\in F(r)$.
Let $\widehat{w}_{2}\in\Sigma_{k}^{*}$ be such that $\widehat{w}_{2}x\in\operatorname{Prf}(\widetilde{w}_{2})$ and $x\not\in F(\widehat{w}_{2})$. Clearly $\widehat{w}_{2}$ exists and is uniquely determined; realize that $\widehat{w}_{2}$ may be the empty word.
Since $p\widehat{w}_{2}x\in F(r^{\beta})$, $x\not\in F(p\widehat{w}_{2})$, and $|p|>|\widetilde{w}_{2}|$ we have that
$p\widehat{w}_{2}x\in F(rr)\mbox{.}$ (2)
Recall $x\not\in F(s_{1})\cup F(s_{2})\cup F(t_{1})\cup F(t_{2})$. Then it follows from (1) that
$\operatorname{occur}(s_{1}\widetilde{w}_{2}hpw_{2}t_{2},p\widehat{w}_{2}x)\leq 2\mbox{.}$ (3)
From (2), (3), and $\beta>\alpha\geq 4$ we conclude that there are no such $r,g,\beta$ and hence $s_{1}x\widetilde{w}_{1}xhpx\widetilde{w}_{2}xt_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
This completes the proof. ∎
The main theorem of the article says that if $w_{1},w_{2}$ are finite nonempty factors of bi-infinite $\alpha$-power free words then there is a bi-infinite $\alpha$-power free word $v$ and a finite word $w$ such that $w_{1}ww_{2}$ is a factor of $v$.
Theorem 4.4.
If $v_{1},v_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $w_{1}\in F(v_{1})\setminus\{\epsilon\}$, and $w_{2}\in F(v_{2})\setminus\{\epsilon\}$ then there are $w\in\Sigma_{k}^{*}$ and $v\in L_{k,\alpha}^{\mathbb{Z}}$ such that $w_{1}ww_{2}\in F(v)$.
Proof.
If $F(v_{1})\cap F(v_{2})=\emptyset$ then let $v_{11}$, $v_{12}$, $v_{21}$, and $v_{22}$ be such that $v_{1}=v_{11}v_{12}$, $v_{2}=v_{21}v_{22}$, $w_{1}\in F(v_{11})$, and $w_{2}\in F(v_{22})$. Obviously $v_{11}{v_{22}}$ is $\alpha$-power free since $F(v_{11})\cap F(v_{22})=\emptyset$.
Thus suppose that $F(v_{1})\cap F(v_{2})\not=\emptyset$ and let $x\in F(v_{1})\cap F(v_{2})\cap\Sigma_{k}$. Without loss of generality we may suppose that $x\in F(w_{1})\cap F(w_{2})$.
Lemma 4.2 asserts that there are $s_{1}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma_{1},\eta_{1}\in\Sigma_{k}^{*}$ such that $(s_{1},\sigma_{1},w_{1},\eta_{1},x)\in\Delta$. Lemma 4.2 asserts also that there are $t_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\sigma_{2},\eta_{2}\in\Sigma_{k}^{*}$ such that $(t_{2}^{R},\sigma_{2}^{R},w_{2}^{R},\eta_{2}^{R},x)\in\Delta$.
Let $t_{1}$ be a right infinite $\alpha$-power free word on the alphabet $\Sigma_{k}\setminus\{x\}$. Since $\alpha\geq 4$, Lemma 2.3 asserts that $t$ exists. Let $s_{2}\in\Phi(t)$; Lemma 2.4 implies that $\Phi(t_{1})\not=\emptyset$. Clearly $s_{2}$ is $\alpha$-power free and $x\not\in F(s_{2})$.
Theorem 4.1 implies that are $\overline{\eta}_{1}\in\operatorname{Prf}(\eta_{1})$ and $\overline{\eta}_{2}\in\operatorname{Suf}(\eta_{2})$ such that $s_{1}\sigma_{1}w_{1}\overline{\eta}_{1}xt_{1},t_{2}^{R}\sigma_{2}^{R}w_{2}^{R}\overline{\eta}_{2}^{R}xs_{2}^{R}\in L_{k,\alpha}^{\mathbb{Z}}$. Since the reverse operation preserves the power-freeness we have also that $s_{2}x\overline{\eta}_{2}w_{2}\sigma_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
Let $\widetilde{w}_{1}=\sigma_{1}w_{1}\overline{\eta}_{1}x$ and let $\widetilde{w}_{2}=x\overline{\eta}_{2}w_{2}\sigma_{2}$. Then Proposition 4.3 implies that there is $w\in\Sigma_{k}^{*}$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$. The theorem follows.
Table 1 shows the structure of the resulting word using also the notation from the proof of Proposition 4.3.
$\overline{w}_{1}$ $\overline{w}_{2}$
$s_{1}$ $\sigma_{1}$ $w_{1}$ $\overline{\eta}_{1}$ $x$ $h$ $p$ $x$ $\overline{\eta}_{2}$ $w_{2}$ $\sigma_{2}$ $t_{2}$
Table 1: The structure of the transition word $v$.
This completes the proof. ∎

### v4

@@
 1 Introduction
-Let $r$ be a non-empty word, let $\alpha$ be a positive rational constant, and let $r^{\alpha}=rr\cdots rt$ be such that $\frac{|r^{\alpha}|}{|r|}=\alpha$ and $t$ is a prefix of $r$.
+Due to a significant overlap of notation, we copy and use the definitions and some explications from sections “Introduction” and “Preliminaries” from [6].
+Let $r$ be a non-empty word, let $\alpha$ be a positive rational number, and let $r^{\alpha}=rr\cdots rt$ be such that $\frac{|r^{\alpha}|}{|r|}=\alpha$ and $t$ is a prefix of $r$.
 We call the word $r^{\alpha}$ an $\alpha$-power of $r$.
 For example $(1234)^{3}=123412341234$, $(1234)^{\frac{7}{4}}=1234123$, and $(1234)^{\frac{1}{2}}=12$.
-The square free ($2$-power free) and cube free words ($3$-power free) are well known examples of the power free words.
-More detailed introduction into the field of power free words can be found, for instance, in [2] and [8].
-In 1985, Restivo and Salemi presented a list of five problems concerning the extendability of power free words [3].
+Suppose a finite or infinite word $w$. Let
+$\begin{split}\Theta(w)=\{(r,\alpha)\mid r^{\alpha}\mbox{ is a factor of }w\mbox{ and }r\mbox{ is a non-empty word and }\\ \alpha\mbox{ is a positive rational number}\}\mbox{.}\end{split}$
+We say that $w$ is $\alpha$-power-free if
+$\{(r,\beta)\in\Theta(w)\mid\beta\geq\alpha\}=\emptyset$
+and we say that $w$ is $\alpha^{+}$-power-free if
+$\{(r,\beta)\in\Theta(w)\mid\beta>\alpha\}=\emptyset\mbox{.}$
+The square free ($2$-power-free) and cube free words ($3$-power-free) are well known examples of the power-free words.
+More detailed introduction into the field of power-free words can be found, for instance, in [3] and [9].
+In 1985, Restivo and Salemi presented a list of five problems concerning the extendability of power-free words [4].
 Problem $4$ and Problem $5$ are relevant for the current article, hence we recall these two problems:
-  • Problem $4$: Given finite $\alpha$-power-free words $u$ and $v$, decide whether there is a transition word $w$, such that $uwu$ is $\alpha$-power free.
+  • Problem $4$: Given finite $\alpha$-power-free words $u$ and $v$, decide whether there is a transition word $w$, such that $uwu$ is $\alpha$-power-free.
   • Problem $5$: Given finite $\alpha$-power-free words $u$ and $v$, find a transition word $w$, if it exists.
-In 2019, a survey on the solution of all the five problems has been presented in [1]; specifically Problem $4$ and Problem $5$ were resolved for some binary language.
-In addition, in [1] the authors resolved Problem $4$ and $5$ for cube free words.
+In 2019, a survey on the solution of all the five problems has been presented in [2]; in particular, Problem $4$ and Problem $5$ were resolved for some binary languages.
+In addition, in [2] the authors resolved Problem $4$ and Problem $5$ for cube free words.
 Let $\mathbb{N}$ denote the set of positive integers and let $\mathbb{Q}$ denote the set of rational numbers.
 Definition 1.1.
-(see [4, Definition $1$])
+([5, Definition $1$]) Let
 $\begin{split}\Upsilon=\{(k,\alpha)\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k=3\mbox{ and }\alpha>2\}\\ \cup\{(k,\alpha)\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k>3\mbox{ and }\alpha\geq 2\}\\ \cup\{(k,\alpha^{+})\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k\geq 3\mbox{ and }\alpha\geq 2\}\mbox{.}\end{split}$
 Remark 1.2.
-(see [4, Remark $1$]) The definition of $\Upsilon$ says that: If $(k,\alpha)\in\Upsilon$ and $\alpha$ is a “number with $+$” then $k\geq 3$ and $\alpha\geq 2$.
-If $(k,\alpha)\in\Upsilon$ and $\alpha$ is “just” a number then $k=3$ and $\alpha>2$ or $k>3$ and $\alpha\geq 2$.
-In 2020, Problem $4$ and Problem $5$ have were resolved for $\alpha$-power free languages over an alphabet with $k$ letters, where $(k,\alpha)\in\Upsilon$ [4].
+([5, Remark $1$]) The definition of $\Upsilon$ says that: If $(k,\alpha)\in\Upsilon$ and $\alpha$ is a number with $+$ then $k\geq 3$ and $\alpha\geq 2$.
+If $(k,\alpha)\in\Upsilon$ and $\alpha$ is just a number then $k=3$ and $\alpha>2$ or $k>3$ and $\alpha\geq 2$.
+In 2020, Problem $4$ and Problem $5$ were resolved for $\alpha$-power-free languages over an alphabet with $k$ letters, where $(k,\alpha)\in\Upsilon$ [5].
 In 2009, a conjecture related to Problems $4$ and Problem $5$ of Restivo and Salemi appeared in [8]:
@@
 In 2018, Conjecture 1.3 appeared also in [7] using a “Restivo Salemi property”; it was defined that a language $L$ has the Restivo Salemi property if Conjecture 1.3 holds for the language $L$.
-Let $L_{k,\alpha}$ denote an $\alpha$-power free language over an alphabet with $k$ letters, where $\alpha$ is a positive rational number and $k$ is positive integer.
-In [6, Remark $41$], it was mentioned that for $2^{+}\leq\alpha\leq\frac{7}{3}$ the language $L_{2,\alpha}$ has the Restivo Salemi property (Conjecture 1.3 holds for $L_{2,\alpha}$), where $2^{+}$-power free language denotes the overlap free language.
+(Thus Restivo-Salemi property is exactly the transition property for extendable power-free languages.)
+[…]
+In [7, Remark $41$], it was mentioned that for $2^{+}\leq\alpha\leq\frac{7}{3}$ the language $L_{2,\alpha}$ has the Restivo Salemi property (Conjecture 1.3 holds for $L_{2,\alpha}$).
+Recall that $2^{+}$-power-free language is the overlap free language.
 Otherwise, as to our best knowledge, Conjecture 1.3 remains open.
@@
 $\widetilde{\Upsilon}=\{(k,\alpha)\mid k\in\mathbb{N}\mbox{ and }\alpha\in\mathbb{Q}\mbox{ and }k\geq 3\mbox{ and }\alpha\geq 5\}\mbox{.}$
-Clearly $\widetilde{\Upsilon}\subset\Upsilon$. In the current article we prove that if $(k,\alpha)\in\widetilde{\Upsilon}$ then Conjecture 1.3 holds for the language $L_{k,\alpha}$.
-In order to prove our result we apply the ideas and some results from [4] and [5].
-We explain briefly the very basic idea.
-If $u,v$ are $\alpha$-power free words and $x$ is a letter such that $x$ is a factor of neither $u$ nor $v$, then clearly $uxv$ is square free ($2$-power free).
+Clearly $\widetilde{\Upsilon}\subset\Upsilon$.
+In the current article, we prove that if $(k,\alpha)\in\widetilde{\Upsilon}$ then Conjecture 1.3 holds for the language $L_{k,\alpha}$; see Corollary 3.10.
+The requirement that $\alpha\geq 5$ was necessary in the proof of Proposition $3$ in [6].
+Since we use the results of [6], we require also for the current paper that $\alpha\geq 5$.
+In order to prove our result, we apply some ideas and results from [5] and [6].
+We briefly explain the very basic idea.
+If $u,v$ are square power-free words ($2$-power-free) and $x$ is a letter such that $x$ is a factor of neither $u$ nor $v$, then clearly $uxv$ is square free.
 Just note that there cannot be a factor in $uxv$ which is a square and contains $x$, because $x$ has only one occurrence in $uxv$.
-In [4], the author considers right and left infinite words containing a given factor and a non-recurrent letter $x$.
-Due to the non-recurrence of the letter $x$ and some involved observations of recurrent and non-recurrent factors, the author managed to prove the results.
-In 2023, the idea of right and left infinite power free words with a non-recurrent letter have been extended to bi-infinite $\alpha$-power free words over an alphabet with $k$-letters, where $(\alpha,k)\in\widetilde{\Upsilon}$, [5].
-Moreover, in [5], it was predicted, that these bi-infinite power free words with a non-recurrent letter could help to prove Conjecture 1.3.
+This very basic idea was generalized to a finite number of occurrences of $x$ instead of a one single occurrence of $x$: In [5], the author constructed right and left infinite words containing a given factor and a non-recurrent letter $x$.
+The non-recurrence of the letter $x$ and several involved observations of recurrent and non-recurrent factors allowed the author to resolve Problem $4$ and Problem $5$ of Restivo and Salemi for languages $L_{k,\alpha}$, where $(k,\alpha)\in\Upsilon$.
+In 2023, the construction of right and left infinite power-free words with a non-recurrent letter has been extended to bi-infinite $\alpha$-power-free words over an alphabet with $k$-letters, where $(k,\alpha)\in\widetilde{\Upsilon}$ [6].
+In [6], it was also predicted that these bi-infinite power-free words with a non-recurrent letter could help to prove Conjecture 1.3.
+In the current article we apply indeed these bi-infinite $\alpha$-power-free words with a non-recurrent letter $x$ to prove Conjecture 1.3 for languages $L_{k,\alpha}$, where $(k,\alpha)\in\widetilde{\Upsilon}$.
 2 Preliminaries
-Due to a significant overlap of notation, we copy and use the section “Preliminaries” from [5].
-Let $\Sigma_{k}$ denote an alphabet with $k$ letters.
-Let $\Sigma_{k}^{+}$ denote the set of all nonempty finite words over $\Sigma_{k}$, let $\epsilon$ denote the empty word, let $\Sigma_{k}^{*}=\Sigma_{k}^{+}\cup\{\epsilon\}$, let $\Sigma_{k}^{\mathbb{N},R}$ denote the set of all right infinite words over $\Sigma_{k}$, let $\Sigma_{k}^{\mathbb{N},L}$ denote the set of all left infinite words over $\Sigma_{k}$, and let $\Sigma_{k}^{\mathbb{Z}}$ denote the set of all bi-infinite words over $\Sigma_{k}$.
+Let $\Sigma_{k}^{+}$ denote the set of all non-empty finite words over $\Sigma_{k}$, let $\epsilon$ denote the empty word, let $\Sigma_{k}^{*}=\Sigma_{k}^{+}\cup\{\epsilon\}$, let $\Sigma_{k}^{\mathbb{N},R}$ denote the set of all right infinite words over $\Sigma_{k}$, let $\Sigma_{k}^{\mathbb{N},L}$ denote the set of all left infinite words over $\Sigma_{k}$, and let $\Sigma_{k}^{\mathbb{Z}}$ denote the set of all bi-infinite words over $\Sigma_{k}$.
 Let $\Sigma_{k}^{\infty}=\Sigma_{k}^{\mathbb{N},L}\cup\Sigma_{k}^{\mathbb{N},R}\cup\Sigma_{k}^{\mathbb{Z}}$. We call $w\in\Sigma_{k}^{\infty}$ an infinite word.
-Let $\operatorname{occur}(w,t)$ denote the number of all occurrences of the nonempty factor $t\in\Sigma_{k}^{+}$ in the word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$.
+Let $\operatorname{occur}(w,t)$ denote the number of all occurrences of the non-empty factor $t\in\Sigma_{k}^{+}$ in the word $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\infty}$.
 If $w\in\Sigma_{k}^{\infty}$ and $\operatorname{occur}(w,t)=\infty$, then we call $t$ a recurrent factor in $w$.
@@
 The set $F(w)$ contains the empty word and if $w$ is finite then also $w\in F(w)$.
-Let $F_{r}(w)\subseteq F(w)$ denote the set of all recurrent nonempty factors of $w\in\Sigma_{k}^{\infty}$.
+Let $F_{r}(w)\subseteq F(w)$ denote the set of all recurrent non-empty factors of $w\in\Sigma_{k}^{\infty}$.
 Let $\operatorname{Prf}(w)\subseteq F(w)$ denote the set of all prefixes of $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\mathbb{N},R}$ and let $\operatorname{Suf}(w)\subseteq F(w)$ denote the set of all suffixes of $w\in\Sigma_{k}^{*}\cup\Sigma_{k}^{\mathbb{N},L}$.
 We define that $\epsilon\in\operatorname{Prf}(w)\cap\operatorname{Suf}(w)$ and if $w$ is finite then also $w\in\operatorname{Prf}(w)\cap\operatorname{Suf}(w)$.
-[…]
+Let $L_{k,\alpha}^{\infty}\subseteq\Sigma_{k}^{\infty}$ denote the set of all infinite $\alpha$-power-free words over $\Sigma_{k}$; formally $L_{k,\alpha}^{\infty}=\{w\in\Sigma_{k}^{\infty}\mid F(w)\subseteq L_{k,\alpha}\}$.
+In addition we define $L_{k,\alpha}^{\mathbb{N},R}=L_{k,\alpha}^{\infty}\cap\Sigma_{k}^{\mathbb{N},R}$, $L_{k,\alpha}^{\mathbb{N},L}=L_{k,\alpha}^{\infty}\cap\Sigma_{k}^{\mathbb{N},L}$, and $L_{k,\alpha}^{\mathbb{Z}}=L_{k,\alpha}^{\infty}\cap\Sigma_{k}^{\mathbb{Z}}$; it means the sets of right infinite, left infinite, and bi-infinite $\alpha$-power-free words, respectively.
 We define the reversal $w^{R}$ of a finite or infinite word $w=\Sigma_{k}^{*}\cup\Sigma_{k}^{\mathbb{N},R}\cup\Sigma_{k}^{\mathbb{N},L}$ as follows:
@@
 Remark 2.1.
-It is obvious that the reverse function preserves the power-freeness and that every factor of an $\alpha$-power free word is also $\alpha$-power free.
-The next proposition is a “reformulation” of Corollary $1$ from [4] using only the notation of the current article.
+It is obvious that the reverse function preserves the power-freeness and that every factor of an $\alpha$-power-free word is also $\alpha$-power-free.
+The next proposition is a reformulation of Corollary $1$ from [5] using only the notation of the current article.
 Proposition 2.2.
-(reformulation of [4, Corollary $1$]) If $(k,\alpha)\in\Upsilon$, $v\in L_{k,\alpha}^{\mathbb{N},L}$, $z\in\operatorname{Suf}(v)$, $x\in F_{r}(v)\cap\Sigma_{k}$, $s\in L_{k,\alpha}^{\mathbb{N},L}$, and $x\not\in F(s)$ then there is a finite word $u\in\Sigma_{k}^{*}$ such that $z\in\operatorname{Suf}(su)$ and $su\in L_{k,\alpha}^{\mathbb{N},L}$.
-The following elementary lemma was shown in [4]. For the reader’s convenience we present the lemma with the proof.
+(reformulation of [5, Corollary $1$]) If $(k,\alpha)\in\Upsilon$, $v\in L_{k,\alpha}^{\mathbb{N},L}$, $z\in\operatorname{Suf}(v)$, $x\in F_{r}(v)\cap\Sigma_{k}$, $s\in L_{k,\alpha}^{\mathbb{N},L}$, and $x\not\in F(s)$ then there is a finite word $u\in\Sigma_{k}^{*}$ such that $z\in\operatorname{Suf}(su)$ and $su\in L_{k,\alpha}^{\mathbb{N},L}$.
+Remark 2.3.
+([6, Remark $3$]) Proposition 2.2 says that if $z$ is a finite $\alpha$-power-free word that can be extended to a left infinite $\alpha$-power-free word having a letter $x$ as a recurrent factor and $s$ is a left infinite $\alpha$-power-free word not containing the letter $x$ as a factor, then there is a left infinite $\alpha$-power-free word containing $z$ as a suffix and having only a finite number of occurrences of $x$.
+The next elementary lemma follows easily from famous theorems of Thue [10, 11].
 Lemma 2.4.
-(reformulation [4, Lemma $2$]) If $k\geq 3$ and $\alpha>2$ then $L_{k-1,\alpha}^{\mathbb{N},R}\not=\emptyset$.
-Proof.
-Thue Morse words are well known overlap-free ($2^{+}$-power free) right infinite words on two letters [6].
-Let $x\in\Sigma_{k}$ and let $t\in L_{k,\alpha}^{\mathbb{N},R}$ be a Thue Morse word on two letters such that $x\not\in F(t)$.
-Since $k\geq 3$, we have that such $t$ exists.
-This ends the proof.
-∎
-Given $t\in\Sigma_{k}^{\mathbb{N},R}$, let
-$\Phi(t)=\{s\in\Sigma_{k}^{\mathbb{N},L}\mid F(s)\subseteq F_{r}(t)\}\mbox{.}$
-It means that all factors of $s\in\Phi(t)$ are recurrent factors of $t$. It was shown in [4], that $\Phi(t)$ is nonempty:
+(reformulation of [5, Lemma $2$]) If $k\geq 3$ and $\alpha>2$ then $L_{k-1,\alpha}^{\mathbb{N},R}\not=\emptyset$.
+Given $t\in\Sigma_{k}^{\mathbb{N},R}$, let $\Phi(t)=\{s\in\Sigma_{k}^{\mathbb{N},L}\mid F(s)\subseteq F_{r}(t)\cup\{\epsilon\}\}\mbox{.}$ It means that all non-empty factors of $s\in\Phi(t)$ are recurrent factors of $t$.
+The following lemma is an immediate consequence of König’s lemma [1].
 Lemma 2.5.
-[4, Lemma $4$] If $t\in\Sigma_{k}^{\mathbb{N},R}$ then $\Phi(t)\not=\emptyset$.
+([5, Lemma $4$]) If $k\in\mathbb{N}$ and $t\in\Sigma_{k}^{\mathbb{N},R}$ then $\Phi(t)\not=\emptyset$.
 The main result of [6] can be formulated as follows:
 Theorem 2.6.
-If $\alpha\geq 5$, $k\geq 3$, $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, then there are $\widetilde{v}\in L_{k,\alpha}^{\mathbb{Z}}$ and a letter $x\in\Sigma$ such that $w\in F(\widetilde{v})$, $x\in F(w)$, and $\operatorname{occur}(w,x)<\infty$.
+If $(k,\alpha)\in\widetilde{\Upsilon}$, $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, then there are $\widetilde{v}\in L_{k,\alpha}^{\mathbb{Z}}$ and a letter $x\in\Sigma_{k}$ such that $w\in F(\widetilde{v})$, $x\in F(w)$, and $\operatorname{occur}(w,x)<\infty$.
 Remark 2.7.
-However, we will not apply Theorem 2.5 directly in the current article, hence we will require some additional properties of the prefix and the suffix of the constructed bi-infinite word.
-Instead we will apply [5, Theorem $1$] (see Theorem 4.1 in the current article).
+We will not apply Theorem 2.6 directly, because we will need some additional properties of the constructed bi-infinite word.
+Instead, we will apply [6, Theorem $1$] (see Theorem 3.4 in the current article).
 3 Bi-infinite $\alpha$-power words
-For the rest of the article suppose that $(k,\alpha)\in\widetilde{\Upsilon}$; it means $k\geq 3$ and $\alpha\geq 5$.
-We define two technical sets $\Gamma$ and $\Delta$.
+For the rest of the article suppose that $(k,\alpha)\in\widetilde{\Upsilon}$; it means $k\geq 3$ and $\alpha\geq 5$. We define two technical sets $\Gamma$ and $\Delta$.
 Definition 3.1.
-([5, Definition $2$]) Let $\Gamma$ be a set of triples such that $(w,\eta,u)\in\Gamma$ if and only if
+([6, Definition $2$]) Let $\Gamma$ be a set of triples such that $(w,\eta,u)\in\Gamma$ if and only if
   • $w\in\Sigma_{k}^{+}$, $\eta,u\in\Sigma_{k}^{*}$, and
@@
 Remark 3.2.
-([5, Remark $4$]) The set $\Gamma$ contains triples of finite words $w,\eta,u$ such that $w$ is nonempty, and if $u$ is shorter than $w$, then the word $\eta$ is “sufficiently” longer than $w$.
+([6, Remark $4$]) The set $\Gamma$ contains triples of finite words $w,\eta,u$ such that $w$ is non-empty, and if $u$ is shorter than $w$, then the word $\eta$ is “sufficiently” longer than $w$.
 Definition 3.3.
-([5, Definition $3$]) Let $\Delta$ be a set of $6$-tuples such that $(s,\sigma,w,\eta,x,u)\in\Delta$ if and only if
+([6, Definition $3$]) Let $\Delta$ be a set of $6$-tuples such that
+$(s,\sigma,w,\eta,x,u)\in\Delta$
+if and only if
   1. 1.
@@
     $x\not\in F(s)\cup F(u)$.
-4 Transition words
+In [6], it was shown that if $(s,\sigma,w,\eta,x,\epsilon)\in\Delta$ and $t$ is a right infinite $\alpha$-power-free word with no occurrence of the letter $x$ then there is a bi-infinite $\alpha$-power-free word containing the factor $w$ and having only a finite number of occurrences of $x$:
 Theorem 3.4.
-([5, Theorem $1$]) If $(s,\sigma,w,\eta,x,\epsilon)\in\Delta$, $t\in L_{k,\alpha}^{\mathbb{N},R}$, and $x\not\in F(t)$ then there is $\widehat{\eta}\in\operatorname{Prf}(\eta)$ such that $s\sigma w\widehat{\eta}xt\in L_{k,\alpha}^{\mathbb{Z}}$.
-We show how construct an element of $\Delta$ from a bi-infinite $\alpha$-power free word $v$ with a finite factor $w$.
+([6, Theorem $1$]) If $(s,\sigma,w,\eta,x,\epsilon)\in\Delta$, $t\in L_{k,\alpha}^{\mathbb{N},R}$, and $x\not\in F(t)$ then there is $\widehat{\eta}\in\operatorname{Prf}(\eta)$ such that $s\sigma w\widehat{\eta}xt\in L_{k,\alpha}^{\mathbb{Z}}$.
+Definition 3.5.
+([6, Definition $4$]) Suppose $v\in L_{k,\alpha}^{\mathbb{Z}}$ and $w\in F(v)\setminus\{\epsilon\}$.
+If there are $v_{1}\in L_{k,\alpha}^{\mathbb{N},L}$ and $v_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ such that $v=v_{1}v_{2}$ and $w\in F_{r}(v_{2})$ then we say that $w$ is on-right-side recurrent in $v$.
+Remark 3.6.
+[6, Remark $8$]) Note in Definition 3.5 that no restriction is imposed on the recurrence of $w$ in $v_{1}$. This means that $w$ may also be recurrent in $v_{1}$.
+We show how to construct an element of $\Delta$ from a bi-infinite $\alpha$-power-free word $v$ with a finite factor $w$ containing the letter $x$.
 Lemma 3.7.
-If $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, and $x\in F(w)\cap\Sigma_{k}$ then there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$, $\sigma,\eta\in\Sigma_{k}^{*}$ such that $(\overline{s},\sigma,w,\eta,x,\epsilon)\in\Delta$.
+If $v\in L_{k,\alpha}^{\mathbb{Z}}$, $w\in F(v)\setminus\{\epsilon\}$, and $x\in F(w)\cap\Sigma_{k}$ then there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$, and $\sigma,\eta\in\Sigma_{k}^{*}$ such that $(\overline{s},\sigma,w,\eta,x,\epsilon)\in\Delta$.
 Proof.
@@
   1. 1.
-    There are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\eta\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\eta x\widehat{v}$ and $(w,\eta,\epsilon)\in\Gamma$.
+    If $\operatorname{occur}(v,w)=\infty$ or $x$ is on-right-side recurrent in $v$ then it is easy to see that there are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$, and $\eta\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\eta x\widehat{v}$ and $(w,\eta,\epsilon)\in\Gamma$.
   2. 2.
-    Otherwise.
-Then it follows from Case 1 that in this case there are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\overline{\eta}\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\overline{\eta}\widehat{v}$ such that $w\not\in F(\overline{v})$, $x\not\in F(\widehat{v})$, and $(w,\overline{\eta},\epsilon)\not\in\Gamma$.
-    Let $u\in\operatorname{Prf}(\widehat{v})$ be such that $|u|>|w\overline{\eta}|$.
-We claim that $\overline{v}w\overline{\eta}ux\in L_{k,\alpha}^{\mathbb{N},L}$.
-To get a contradiction, suppose that there are $r$ and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(\overline{v}w\overline{\eta}ux)$.
-Since $x\not\in F(u)$ and $u\not=\epsilon$ we have that $ux\in L_{k,\alpha}$ and $ux\in\operatorname{Suf}(r)$.
-Since $|u|>|w\overline{\eta}|$ it follows that $w\in F(rr)$.
-Since $\beta>\alpha\geq 4$ and $w\not\in F(\overline{v})$, we conclude that are not such $r$ and $\beta$.
-Set $\eta=\overline{\eta}u$.
-Obviously $(w,\eta,\epsilon)\in\Gamma$, because $(w,\overline{\eta},\epsilon)\in\Gamma$.
-In both cases we found $\overline{v}$ and $\eta$ such that $\overline{v}w\eta x\in L_{k,\alpha}^{\mathbb{N},L}$ and $(w,\eta,\epsilon)\in\Gamma$.
-  • If $x\not\in F(\overline{v})$ then let $\overline{s}=\overline{v}$. It is easy to verify that $(\overline{s},\sigma,w,\eta,x)\in\Delta$.
-  • If $x\in F(\overline{v})$ then let $t$ be a right infinite $\alpha$-power free word on the alphabet $\Sigma_{k}\setminus\{x\}$.
-Since $(\alpha,k)\in\widetilde{\Upsilon}$, Lemma 2.3 asserts that $t$ exists.
+    Otherwise; it means that $\operatorname{occur}(v,w)<\infty$ and $x$ is not on-right-side recurrent in $v$.
+Then there are $\overline{v}\in L_{k,\alpha}^{\mathbb{N},L}$, $\widehat{v}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\overline{\eta}\in\Sigma_{k}^{*}$ such that $v=\overline{v}w\overline{\eta}\widehat{v}$,
+    $x\not\in F(\widehat{v})\mbox{, }$ (1)
+    and
+    $\operatorname{occur}(\overline{v}w,w)=1\mbox{.}$ (2)
+    Let $\eta_{0}\in\operatorname{Prf}(\widehat{v})$ be such that
+    $|\eta_{0}|>\max\{|w\overline{\eta}|,(\alpha+1)\alpha^{|w|}|w|\}\mbox{.}$ (3)
+    We claim that $\overline{v}w\overline{\eta}\eta_{0}x\in L_{k,\alpha}^{\mathbb{N},L}$.
+To get a contradiction, suppose that there are $r$ and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(\overline{v}w\overline{\eta}\eta_{0}x)$.
+From (1) we have that $x\not\in F(\eta_{0})$; consequently $\eta_{0}x\in L_{k,\alpha}$ and $\eta_{0}x\in\operatorname{Suf}(r)$.
+From (3) we have that $|\eta_{0}|>|w\overline{\eta}|$; it follows that $w\in F(rr)$.
+From $\beta>\alpha\geq 5$ and (2), we conclude that are not such $r$ and $\beta$.
+Set $\eta=\overline{\eta}\eta_{0}$.
+From (3) it is obvious $(w,\eta,\epsilon)\in\Gamma$.
+In both cases we found $\overline{v}$ and $\eta$ such that $\overline{v}w\eta x\in L_{k,\alpha}^{\mathbb{N},L}$, and $(w,\eta,\epsilon)\in\Gamma$.
+  • If $x\not\in F(\overline{v})$ then let $\overline{s}=\overline{v}$. It is obvious that $(\overline{s},\epsilon,w,\eta,x,\epsilon)\in\Delta$.
+  • If $x\in F(\overline{v})$ then let $t$ be a right infinite $\alpha$-power-free word on the alphabet $\Sigma_{k}\setminus\{x\}$.
+Since $(k,\alpha)\in\widetilde{\Upsilon}$, Lemma 2.4 asserts that $t$ exists.
 Let $s=t^{R}$.
-Clearly $s$ is $\alpha$-power free and $x\not\in F(s)$.
+Clearly $s$ is a left infinite $\alpha$-power-free word and $x\not\in F(s)$.
     Let $z=w\eta x$.
-Proposition 2.2 implies that there is a finite word $u$ such that $su\in L_{k,\alpha}^{\mathbb{N},L}$ and $z=w\eta x\in\operatorname{Suf}(su)$.
-Obviously there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma\in\Sigma_{k}^{*}$ such that $(\overline{s},\sigma,w,\eta,x)\in\Delta$.
-Note that since $x\in F(w)$ and $x\not\in F(\overline{s})$ Property 4 of Definition 3.3 is easy to be asserted by a proper choice of $\sigma$.
+Proposition 2.2 implies that there is a finite word $u$ such that $su\in L_{k,\alpha}^{\mathbb{N},L}$ and $z\in\operatorname{Suf}(su)$.
+Obviously there are $\overline{s}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma,\eta_{0}\in\Sigma_{k}^{*}$ such that $\overline{s}\sigma w\eta_{0}x=su$, $\eta\in\operatorname{Suf}(\eta_{0})$, $w\eta x\in\operatorname{Suf}(\overline{s}\sigma w\eta_{0}x)$, and $(\overline{s},\sigma,w,\eta_{0},x,\epsilon)\in\Delta$.
+Note that since $x\in F(w)$ and $x\not\in F(\overline{s})$ Property 4 of Definition 3.3 is easy to be asserted by a proper choice of $\sigma$ and $\eta_{0}$.
 This completes the proof.
-Suppose two bi-infinite $\alpha$-power free words $s_{1}\widetilde{w}_{1}t_{1}$ and $s_{2}\widetilde{w}_{2}t_{2}$, where $s_{1},s_{2}$ are left infinite words, $t_{1},t_{2}$ are right infinite words, $w_{1},w_{2}$ are finite words, the factors of $s_{2}$ are recurrent factors of $t_{1}$, and there is a letter $x$ such that $x$ is not a factor of $s_{1},s_{2},t_{1},t_{2}$ and $x$ is a factor of $w_{2}$.
-We show that there is a finite word $w$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}$ is a bi-infinite $\alpha$-power free word.
+Suppose two bi-infinite $\alpha$-power-free words $s_{1}\widetilde{w}_{1}t_{1}$ and $s_{2}\widetilde{w}_{2}t_{2}$, where $s_{1},s_{2}$ are left infinite words, $t_{1},t_{2}$ are right infinite words, $\widetilde{w}_{1},\widetilde{w}_{2}$ are finite words, the factors of $s_{2}$ are recurrent factors of $t_{1}$, and there is a letter $x$ such that $x$ is not a factor of $s_{1},s_{2},t_{1},t_{2}$ and $x$ is a factor of $\widetilde{w}_{2}$.
+We show that there is a finite word $w$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}$ is a bi-infinite $\alpha$-power-free word.
 Proposition 3.8.
-If $\widetilde{w}_{1},\widetilde{w}_{2}\in\Sigma_{k}^{+}$, $s_{1},s_{2}\in\Sigma_{k}^{\mathbb{N},L}$, $t_{1},t_{2}\in\Sigma_{k}^{\mathbb{N},R}$, $x\in\Sigma_{k}$, $s_{1}\widetilde{w}_{1}t_{1},s_{2}\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $F(s_{2})\subseteq F_{r}(t_{1})$, $x\not\in F(s_{1})\cup F(s_{2})\cup F(t_{1})\cup F(t_{2})$, and $x\in F(\widetilde{w}_{2})$ then there is $w\in\operatorname{Prf}(t_{1})$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
+If $\widetilde{w}_{1},\widetilde{w}_{2}\in\Sigma_{k}^{+}$, $s_{1},s_{2}\in\Sigma_{k}^{\mathbb{N},L}$, $t_{1},t_{2}\in\Sigma_{k}^{\mathbb{N},R}$, $x\in\Sigma_{k}$, $s_{1}\widetilde{w}_{1}t_{1},s_{2}\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $F(s_{2})\subseteq F_{r}(t_{1})$, $x\not\in F(s_{1})\cup F(s_{2})\cup F(t_{1})\cup F(t_{2})$, and $x\in\operatorname{Prf}(\widetilde{w}_{2})$ then there is $w\in\operatorname{Prf}(t_{1})$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
 Proof.
-Let $p\in\operatorname{Suf}(s_{2})$ be the shortest suffix such that
+Let $p\in\operatorname{Suf}(s_{2})$ be the shortest suffix of $s_{2}$ such that
 $|p|>\max\{|\widetilde{w}_{1}|,|\widetilde{w}_{2}|\}\mbox{.}$ (4)
@@
 We show that $s_{1}\widetilde{w}_{1}hp\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
-To get a contradiction, suppose that there are $g\in\operatorname{Prf}(\widetilde{w}_{2}xt_{2})$, $y\in\Sigma_{k}$, $r\in\Sigma_{k}^{+}$, and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(s_{1}\widetilde{w}_{1}hpgy)$, $gy\in\operatorname{Prf}(\widetilde{w}_{2}xt_{2})$, and $s_{1}\widetilde{w}_{1}hpg\in L_{k,\alpha}^{\mathbb{N},L}$.
-Since $p\widetilde{w}_{2}xt_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $x\in F(w_{2})$ we have that $|r^{\beta}|>|pgy|$ and $x\in F(r)$.
+To get a contradiction, suppose that there are $g\in\operatorname{Prf}(\widetilde{w}_{2}t_{2})$, $y\in\Sigma_{k}$, $r\in\Sigma_{k}^{+}$, and $\beta>\alpha$ such that $r^{\beta}\in\operatorname{Suf}(s_{1}\widetilde{w}_{1}hpgy)$, $gy\in\operatorname{Prf}(\widetilde{w}_{2}t_{2})$, and $s_{1}\widetilde{w}_{1}hpg\in L_{k,\alpha}^{\mathbb{N},L}$.
+Since $p\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $x\in\operatorname{Prf}(\widetilde{w}_{2})$ we have that $|r^{\beta}|>|pgy|$ and $x\in F(r)$.
 Let $\widehat{w}_{2}\in\Sigma_{k}^{*}$ be such that $\widehat{w}_{2}x\in\operatorname{Prf}(\widetilde{w}_{2})$ and $x\not\in F(\widehat{w}_{2})$.
 Clearly $\widehat{w}_{2}$ exists and is uniquely determined; realize that $\widehat{w}_{2}$ may be the empty word.
-Since $p\widehat{w}_{2}x\in F(r^{\beta})$, $x\not\in F(p\widehat{w}_{2})$, and $|p|>|\widetilde{w}_{2}|$ we have that
+Since $p\widehat{w}_{2}x\in F(r^{\beta})$ and $x\not\in F(p\widehat{w}_{2})$, we have that
 $p\widehat{w}_{2}x\in F(rr)\mbox{.}$ (5)
-Recall $x\not\in F(s_{1})\cup F(s_{2})\cup F(t_{1})\cup F(t_{2})$. Then it follows from (1) that
-$\operatorname{occur}(s_{1}\widetilde{w}_{2}hpw_{2}t_{2},p\widehat{w}_{2}x)\leq 2\mbox{.}$ (3)
-From (2), (3), and $\beta>\alpha\geq 4$ we conclude that there are no such $r,g,\beta$ and hence $s_{1}x\widetilde{w}_{1}xhpx\widetilde{w}_{2}xt_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
+Recall $x\not\in F(s_{1})\cup F(hp)\cup F(t_{2})$. Then it follows from (4) that
+$\operatorname{occur}(s_{1}\widetilde{w}_{1}hp\widetilde{w}_{2}t_{2},p\widehat{w}_{2}x)\leq 2\mbox{.}$ (6)
+From (5), (6), and $\beta>\alpha\geq 5$ we conclude that there are no such $r,g,\beta$ and hence $s_{1}\widetilde{w}_{1}hp\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
 This completes the proof. ∎
-The main theorem of the article says that if $w_{1},w_{2}$ are finite nonempty factors of bi-infinite $\alpha$-power free words then there is a bi-infinite $\alpha$-power free word $v$ and a finite word $w$ such that $w_{1}ww_{2}$ is a factor of $v$.
+The main theorem of the article says that if $w_{3},w_{4}$ are finite non-empty factors of bi-infinite $\alpha$-power-free words then there is a bi-infinite $\alpha$-power-free word $v$ and a finite word $w_{0}$ such that $w_{3}w_{0}w_{4}$ is a factor of $v$.
 Theorem 3.9.
-If $v_{1},v_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $w_{1}\in F(v_{1})\setminus\{\epsilon\}$, and $w_{2}\in F(v_{2})\setminus\{\epsilon\}$ then there are $w\in\Sigma_{k}^{*}$ and $v\in L_{k,\alpha}^{\mathbb{Z}}$ such that $w_{1}ww_{2}\in F(v)$.
+If $v_{1},v_{2}\in L_{k,\alpha}^{\mathbb{Z}}$, $w_{3}\in F(v_{1})\setminus\{\epsilon\}$, and $w_{4}\in F(v_{2})\setminus\{\epsilon\}$ then there are $w_{0}\in\Sigma_{k}^{*}$ and $v\in L_{k,\alpha}^{\mathbb{Z}}$ such that $w_{3}w_{0}w_{4}\in F(v)$.
 Proof.
-If $F(v_{1})\cap F(v_{2})=\emptyset$ then let $v_{11}$, $v_{12}$, $v_{21}$, and $v_{22}$ be such that $v_{1}=v_{11}v_{12}$, $v_{2}=v_{21}v_{22}$, $w_{1}\in F(v_{11})$, and $w_{2}\in F(v_{22})$.
-Obviously $v_{11}{v_{22}}$ is $\alpha$-power free since $F(v_{11})\cap F(v_{22})=\emptyset$.
-Thus suppose that $F(v_{1})\cap F(v_{2})\not=\emptyset$ and let $x\in F(v_{1})\cap F(v_{2})\cap\Sigma_{k}$. Without loss of generality we may suppose that $x\in F(w_{1})\cap F(w_{2})$.
-Lemma 4.2 asserts that there are $s_{1}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma_{1},\eta_{1}\in\Sigma_{k}^{*}$ such that $(s_{1},\sigma_{1},w_{1},\eta_{1},x)\in\Delta$.
-Lemma 4.2 asserts also that there are $t_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\sigma_{2},\eta_{2}\in\Sigma_{k}^{*}$ such that $(t_{2}^{R},\sigma_{2}^{R},w_{2}^{R},\eta_{2}^{R},x)\in\Delta$.
-Let $t_{1}$ be a right infinite $\alpha$-power free word on the alphabet $\Sigma_{k}\setminus\{x\}$.
-Since $\alpha\geq 4$, Lemma 2.3 asserts that $t$ exists.
-Let $s_{2}\in\Phi(t)$; Lemma 2.4 implies that $\Phi(t_{1})\not=\emptyset$.
-Clearly $s_{2}$ is $\alpha$-power free and $x\not\in F(s_{2})$.
-Theorem 4.1 implies that are $\overline{\eta}_{1}\in\operatorname{Prf}(\eta_{1})$ and $\overline{\eta}_{2}\in\operatorname{Suf}(\eta_{2})$ such that $s_{1}\sigma_{1}w_{1}\overline{\eta}_{1}xt_{1},t_{2}^{R}\sigma_{2}^{R}w_{2}^{R}\overline{\eta}_{2}^{R}xs_{2}^{R}\in L_{k,\alpha}^{\mathbb{Z}}$.
-Since the reverse operation preserves the power-freeness we have also that $s_{2}x\overline{\eta}_{2}w_{2}\sigma_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
-Let $\widetilde{w}_{1}=\sigma_{1}w_{1}\overline{\eta}_{1}x$ and let $\widetilde{w}_{2}=x\overline{\eta}_{2}w_{2}\sigma_{2}$.
+If $F(v_{1})\cap F(v_{2})=\emptyset$ then let $v_{11}$, $v_{12}$, $v_{21}$, and $v_{22}$ be such that $v_{1}=v_{11}v_{12}$, $v_{2}=v_{21}v_{22}$, $w_{3}\in F(v_{11})$, and $w_{4}\in F(v_{22})$.
+It is clear $v_{11}{v_{22}}$ is $\alpha$-power-free since $F(v_{11})\cap F(v_{22})=\emptyset$.
+The theorem follows.
+Thus suppose that $F(v_{1})\cap F(v_{2})\not=\emptyset$ and let $x\in F(v_{1})\cap F(v_{2})\cap\Sigma_{k}$.
+Let $w_{1}\in F(v_{1})$ and $w_{2}\in F(v_{2})$ be such that $x\in F(w_{1})\cap F(w_{2})$, $w_{3}\in F(w_{1})$, and $w_{4}\in F(w_{2})$.
+It is clear that such $w_{1},w_{2}$ exist.
+Lemma 3.7 asserts that there are $s_{1}\in L_{k,\alpha}^{\mathbb{N},L}$ and $\sigma_{1},\eta_{1}\in\Sigma_{k}^{*}$ such that $(s_{1},\sigma_{1},w_{1},\eta_{1},x,\epsilon)\in\Delta$.
+Lemma 3.7 asserts also that there are $t_{2}\in L_{k,\alpha}^{\mathbb{N},R}$ and $\sigma_{2},\eta_{2}\in\Sigma_{k}^{*}$ such that $(t_{2}^{R},\sigma_{2}^{R},w_{2}^{R},\eta_{2}^{R},x,\epsilon)\in\Delta$.
+Let $t_{1}\in L_{k-1,\alpha}^{\mathbb{N},R}$ be a right infinite $\alpha$-power-free word on the alphabet $\Sigma_{k}\setminus\{x\}$.
+Since $\alpha\geq 5$, Lemma 2.4 asserts that $t_{1}$ exists.
+Let $s_{2}\in\Phi(t_{1})$; Lemma 2.5 implies that $\Phi(t_{1})\not=\emptyset$ and consequently $s_{2}$ exists.
+Clearly $s_{2}$ is $\alpha$-power-free and $x\not\in F(s_{2})$.
+Theorem 3.4 implies that there are $\widehat{\eta}_{1}\in\operatorname{Prf}(\eta_{1})$ and $\widehat{\eta}_{2}\in\operatorname{Suf}(\eta_{2})$ such that $s_{1}\sigma_{1}w_{1}\widehat{\eta}_{1}xt_{1},t_{2}^{R}\sigma_{2}^{R}w_{2}^{R}\widehat{\eta}_{2}^{R}xs_{2}^{R}\in L_{k,\alpha}^{\mathbb{Z}}$.
+Since the reverse operation preserves the power-freeness we have also that $s_{2}x\widehat{\eta}_{2}w_{2}\sigma_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
+Let $\widetilde{w}_{1}=\sigma_{1}w_{1}\widehat{\eta}_{1}x$ and let $\widetilde{w}_{2}=x\widehat{\eta}_{2}w_{2}\sigma_{2}$.
 Then Proposition 3.8 implies that there is $w\in\Sigma_{k}^{*}$ such that $s_{1}\widetilde{w}_{1}w\widetilde{w}_{2}t_{2}\in L_{k,\alpha}^{\mathbb{Z}}$.
-The theorem follows.
 Table 1 shows the structure of the resulting word using also the notation from the proof of Proposition 3.8.
-$\overline{w}_{1}$ $\overline{w}_{2}$
-$s_{1}$ $\sigma_{1}$ $w_{1}$ $\overline{\eta}_{1}$ $x$ $h$ $p$ $x$ $\overline{\eta}_{2}$ $w_{2}$ $\sigma_{2}$ $t_{2}$
-Table 1: The structure of the transition word $v$.
-This completes the proof. ∎
+$\widetilde{w}_{1}$ $\widetilde{w}_{2}$
+$s_{1}$ $\sigma_{1}$ $w_{1}$ $\widehat{\eta}_{1}$ $x$ $h$ $p$ $x$ $\widehat{\eta}_{2}$ $w_{2}$ $\sigma_{2}$ $t_{2}$
+Table 1: The structure of the word $v$.
+Since $w_{3}\in F(w_{1})$, $w_{4}\in F(w_{2})$, $w_{1}\in F(\widetilde{w}_{1})$, and $w_{2}\in F(\widetilde{w}_{2})$, the theorem follows. This completes the proof. ∎
+Theorem 3.9 has the following obvious corollary.
+Corollary 3.10.
+Conjecture 1.3 holds for the languages $L_{k,\alpha}$, where $\alpha\geq 5$ and $k\geq 3$.